evaluate the function g(x)= 1+2x
g(a+1) =

Answer by jim_thompson5910(35256)

g(x)= 1+2x

g(a+1)= 1+2(a+1)

g(a+1)= 1+2a+2

g(a+1)= 2a+3


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim